As of the end of last year there was 6,600,000 square feet occupied in the market, out of an available supply of 7,300,000. During this year there was new construction of 50,000 sq. ft., and an absorption of 150,000 sq. ft. Given these figures, what is the occupancy rate as of the end of this year?

91.84%

93.15%

90.41%

88.00%
Answer #1

A. 91.84%.

total supply available = 7,300,000.+50,000 new construction

=>7,350,000

total occupancy =6,600,000+150,000

=>6,750,000.

occupancy rate = 6,750,000 / 7,350,000 *100

=>91.84%.